OpenStudy (anonymous):

(3x10^-6)(6x10-^8) scientific notation

sam (.sam.):

(3*10^(-6))(6*10^(8)) (3*1/(1000000)) (6*10^(8)) ((3)/(1000000))(6*10^(8)) ((3)/(1000000))(6*100000000) ((3)/(1000000))(600000000) 1800

OpenStudy (anonymous):

I'm assuming that that is supposed to be 10^-8. \[ 18 \times 10^{-14}\]

OpenStudy (anonymous):

thats not eithe of my options the 18 is a 1.8 but the -14 is off

OpenStudy (anonymous):

Is the second term supposed to be \[10^{-8}\]?

OpenStudy (anonymous):

its between -13and -16 i do have an option for 18 is 18 correct for sure

OpenStudy (anonymous):

If so then 3 x 10^-6 = 0.000003 and 6 x 10^-8 = 0.00000006

OpenStudy (anonymous):

Uh, -14 is between -13 & -16. /sigh

OpenStudy (anonymous):

i saying one of those is either to the -13 or -15 or -16 which one

OpenStudy (anonymous):

Assume x = 10. Then you have the terms \[3 \times x^{-6} \times 6 \times x^{-8} = 18 \times x^{-6 + (-8)} = 18 \times x^{-14}\] Then substitute back the 10.

OpenStudy (anonymous):

Oh, I'm silly. True scientific notation should be in the form a.b x 10^c where a, b and c are single digit integers. Thus 18 x 10^-14 = 1.8 x 10 x 10^-14 = 1.8 x 10 ^ -13